Please help me with these:
1. If 7 men can do 3 jobs in 27 hours then ? men can do 5 jobs in 35 hours.
2. If 10 men can do 6 jobs in 48 hours then 6 men can do ? jobs in 40 hours.
3. If 8 men can do 5 jobs in 10 hours then 5 men can do 10 jobs in ? hours.
Thank you.

Answer by Edwin McCravy(20054)

1.  If 7 men can do 3 jobs in 27 hours then ? men can do 5 jobs in 35 hours.

The easiest way to do these problems is to use the fact that

k = (Number of workers) × (Number of hours*) ÷ (Number of jobs)

*(or minutes or days in some problems)

where k is a constant for any number of workers, hours or jobs,
So we calculate the constant k from the first situation:

7 men can do 3 jobs in 27 hours

7×27÷3 = 63
 
Then we let the answer to the problem be 'N men', and we have
for the second situation:

N men can do 5 jobs in 35 hours

So

N×35÷5 must equal the constant k = 63

N×35÷5 = 63

   N×7 = 63 

     N = 63÷7

     N = 9

It would take 9 men.

Edwin